Write the equation in slope-intercept form of the line that is parallel to the graph of each equation and passes through the given point.
!. y = 3x + 6; (4,7)

Answer by VirtualMathTutor(26)

When two lines are parallel, their slopes are equal. Since the slope of
y = 3x + 6
y = mx + b
is 3, then our slope is also m = 3
To find the equation, use the formula y - y1 = m(x - x1)
x1 = 4
y1 = 7
m = 3
y - 7 = 3(x - 4)
y - 7 = 3x - 12
Add 7 to both sides:
y = 3x - 12 + 7
y = 3x - 5